# BatchLLM: Optimizing Large Batched LLM Inference with Global Prefix Sharing and Throughput-oriented Token Batching

### v2

@@ 5. Implementation
Figure 5. Microbenchmark evaluation. The setting m/n (like 2000/200) indicates the length of shared prefix/non-shared context, sd means sharing degree. The vLLM setting with ’+ p’ (’+ c’) means prefix-caching (chunked-prefill) enabled.
6. Evaluation
6.1. Setup
@@
Baseline Specification. We compare BatchLLM with vLLM and SGLang, two state-of-the-art LLM inference engines with prefix sharing, for the end-to-end comparison to demonstrate the efficacy (Sec.6.2). We integrate all our design in vLLM v0.6.4 and the baseline for comparison is also vLLM v0.6.4 (except as noted). We use FlashAttention (Dao et al., 2022) as the Attention backend888 We also tested FlashInfer (Ye et al., 2025) as the backend of vLLM v0.6.4, and found that the gap between FlashAttention and FlashInfer was quite small, and the end-to-end results of FlashAttention were slightly better than the other one. of the vLLM baseline in NVIDIA GPU environments. The version of SGLang is v0.4.1.
@@
We enable the prefix-caching and chunked-prefill in vLLM baseline for the end-to-end comparison in Sec.6.2. The prefix-caching uses the implicit LRU-based caching policy. We use 2048 as the token-batch size when enabling the chunked-prefill of vLLM baseline to maximize its throughput.
For the kernel comparison in Sec.6.3.3, we compare BatchLLM with Cascade-Inference (Ye et al., 2024a), which is the SOTA implementation of the prefix-shared Attention on NVIDIA platform. It is worth noting that Cascade-Inference is similar to our design on kernel-levels. One difference between Cascade-Inference and BatchLLM is that BatchLLM applies horizontal fused kernel. Besides, Cascade-Inference cannot be applied on the non-NVIDIA backend as it is developed with CUDA, while BatchLLM is based on OpenAI Triton and can support both NVIDIA and AMD GPUs. All the other packages like PyTorch and Huggingface transformers are dependent to the docker file of vLLM v0.6.4.
@@
Hardware and System Specification. We conduct our experiments on both NVIDIA A100 80GB GPU and AMD MI200 GPU. The CUDA version on the NVIDIA platform is 12.1. The ROCm version on the AMD platform is 6.1. The CPU is AMD EPYC 7V12 CPU @2.45GHz. The operation system is Ubuntu 20.04.
@@
Models and Dataset. We evaluate BatchLLM and the baselines with Llama-3.1-8B (AI@Meta, 2024), Qwen-2.5-14B (Qwen et al., 2025) and Llama-3.1-70B (4-bit GPTQ), with a set of microbenchmarks with different data distributions and Qwen-2.5-7B for a typical industry task. As for the microbenchmarks evaluation, we use three different settings of the length of common prefix and distinct prompt. The length of prefix/distinct are 2000/200, 2000/1000 and 1000/1000 respectively, while the length of generated tokens is 100. We evaluate the three microbenchmarks under different share-degrees. The share-degree means the number of distinct prompts sharing the same prefix. Note that all requests are randomly shuffled. Specifically, we conduct the evaluation of the three microbenchmarks with sharing degree 4 and 16, respectively. The microbenchmark evaluation is in Sec.6.2.1. As for the industry tasks, we conduct the evaluation on a typical scenario in web search engines, which is the snippet generation task that generates the snippet of a document for the user query.
We conduct the breakdown experiments in Sec.6.3 to demonstrate the efficacy of each techniques in this paper.
6.2. End to End Comparison
@@
Figure 6. Microbenchmark evaluation of different sharing-degrees and different shared-prefix lengths on A100. Given 6400 requests. This evaluation is based on Llama-3.1-8B-Instruct, while the length of non-shared context is 200, and the generated length is 100.
6.2.1. Microbenchmark Evaluation
Fig.5 shows the end-to-end throughput comparison between BatchLLM and the baselines with different shared-prefix/non-shared context settings. In these benchmark experiments, we compare BatchLLM with vLLM and SGLang of different configurations, and with or without prefix-caching and chunked-prefill enabled.
The results show that BatchLLM outperforms the vLLM and SGLang baseline under different configurations, for different models on different GPU platforms with different input data distributions. Specifically, when comparing with vLLM + prefix-caching + chunked-prefill, which performs the best across all vLLM’s configurations, BatchLLM can get 1.4-3$\times$ speedup on A100 and 1.8-4.1$\times$ speedup on MI200. BatchLLM also achieves better performance than SGLang, with 1.2-2.5$\times$ speedup on A100 and 1.3-2.8$\times$ speedup on MI200. The acceleration is stable from 7B to 70B models.
@@ 6.2.1. Microbenchmark Evaluation
The impact of sharing-degree and shared prefix. To show the benefit of BatchLLM, we also compare the throughput of different settings after changing the sharing-degree and the length of shared prefix. Fig.6 shows that the speedup becomes more significant when more requests share the same prefix and the portion of the common prefix becomes larger, which contributes to a bigger reuse ratio, and BatchLLM can save more computation and memory for this case. With lower sharing-degree and shorter common prefix (like sd is 2 and the length of prefix is 2000), the acceleration of BatchLLM is 1.7×1.7$\times$1.7 ×/1.5×1.5$\times$1.5 ×, comparing with the best setting of vLLM and SGLang. When the length of shared prefix is 16000, and the share degree is 16, the best performance of vLLM and SGLang are 0.49 and 0.56, while BatchLLM could achieve 10.8×10.8$\times$10.8 ×/9.5×9.5$\times$9.5 × acceleration.
@@
Comparison between vLLM + request sorting and BatchLLM. We perform a preprocess for the setting of vLLM + chunked-prefill + prefix-cahcing: use python.sorted() to gather all requests with the same prefix together999 Here the length of shared prefix/non-shared context/generated tokens are 2000/200/100, while the sharing degree is 16. The model is Llama-3.1-8B. And the GPU is NVIDIA A100.. This can help vLLM to cache the tokens better. After this preprocess, vLLM’s throughput is boosted from 6 to 13.02, while the throughput of BatchLLM is 18. One reason is that vLLM cannot simplify the attention calculation when there are more than 1 common prefix in the batch every step of its inference, while BatchLLM could handle it. Besides, BatchLLM has a better token-batching strategy than vLLM.
@@
Table 1. Industry workload evaluation.
@@
BatchLLM 8.67(1.30×1.30$\times$1.30 ×) 4.24(1.26×1.26$\times$1.26 ×)
@@
Figure 7. The length distribution of shared prefix/non-shared context in the dataset of the industry workload.
6.2.2. Industry Workload Evaluation
@@
This industry workload is similar to the snippet generation task described in the recent work (Wang et al., 2023), which is also described in Sec.3.1. The original input has two levels of prefixes: the global shared instruction to extract the information from the document-query pair, and the shared document between different groups of queries. The first-level prefix is very short. With the optimization described in Sec.4.2, the first-level prefixes are enlarged (i.e., some of the second-level prefix of the document is merged into the first-level prefix). We have analyzed the token saving ratio according to Eq.2. It shows that the saving ratio are nearly the same between the multi-level and the enlarged single level prefixes, 56% for the former and 55% for the latter.
@@
We synthesize the dataset for the evaluation according to the data distribution of the industry workload (Fig.7), where the average share-degree (i.e., the number of the distinct prompt that share the same prefix) is about 3, the average common prefix length is about 1570 tokens, and the average distinct prompt length is about 30 tokens. We sample 8000 queries for this case study.
We conduct this experiment on the NVIDIA A100 GPU and AMD MI200 GPU, using Qwen-2.5-7B model. Table.1 shows the effectiveness of BatchLLM of this workload (and another industry workload that will describe later), with about 1.3$\times$/1.26$\times$ speedup over the best configuration of vLLM. BatchLLM better reuses the KV context than the vLLM baseline with the optimization in Sec.4.2, for which we have the breakdown analysis in Sec.6.3.1. BatchLLM also better mixes the decoding tokens with prefill tokens to increase the size of token-batches with the optimization in Sec.4.3, for which we have the breakdown analyze in Sec.6.3.2. It also benefits from the optimized Attention kernel described in this paper.
6.3. Breakdown Analysis
6.3.1. Reusing Effect Comparison
@@
To show that BatchLLM optimizes the KV reuse, we evaluate the throughput and the token saving ratio in both microbenchmarks and the industry task.
@@
Figure 8. The token saving (or reusing) ratio of different sharing-degrees and different shared-prefix lengths on A100, based on the microbenchmark. This evaluation is based on Llama-3.1-8B-Instruct, while the length of non-shared context is 200, and the generated length is 100.
@@
As shown in Fig.8, we can see the token saving ratio is improved in BatchLLM across different settings. And it becomes higher when the sharing-degree of requests is larger. Specifically, in the microbenchmark setting where length of shared prefix is 2000 and the sharing-degree is 16, BatchLLM achieves 85.2% token saving ratio, the upper bound in this scenario, comparing to the 40.1% token saving ratio in vLLM + chunked-prefill + prefix-caching. For the industry workload, BatchLLM also makes a token saving ratio of 58.1%, while the number of vLLM + chunked-prefill + prefix-caching is 35.8%. It is worth mentioning that BatchLLM can achieve the best token reusing since it performs explicit common prefix management. With the increase of the requests number and the length of prefix, it can be expected that the implicit LRU cache policy of vLLM’s prefix-caching can suffer more from the eviction of the reusable KV context, while the inference could benefit more from BatchLLM. For example, when the length of shared prefix is 16000, and the share degree is 16, the token reusing ratio of vLLM and SGLang are 6.3% and 5.2%, while the reusing ratio of BatchLLM is 92.6 %, much better than the other two baselines.
6.3.2. Token-batching Analysis
@@
Fig.9 shows the per-iteration token number, using the same workload with Fig.2. It clearly shows that the token-batching optimization in BatchLLM successfully reduces the "valleys" of the iterations, thus can increase the overall GPU utilization.
@@
Figure 9. Token number per iteration with token-batching optimization (Sec.4.3), showing significant valley reduction compared to Fig.2’s baseline. Prefix-sharing disabled to demonstrate general batched/offline inference efficacy.
Table 2. Token-Batching Breakdown.
@@
Figure 10. The performance comparison between the baseline, CascadeInfer and BatchLLM. The setting i/j/k (like 256/2048/all) means the length of global shared prefix, the length of distinct parts for each request, and how many requests with one single shared prefix in one single prefix-sharing group.
6.3.3. Kernel Performance Comparison
We compare BatchLLM’s prefix-shared Attention implementation with several baselines on both NVIDIA A100 GPU and AMD MI200 GPU. Fig.10 shows the performance comparison of different implementations, including Triton official FlashAttention implementation [^10] https://github.com/triton-lang/kernels/blob/main/kernels/flash_attention.py, official FlashAttention [^11] https://github.com/Dao-AILab/flash-attention, v2.6.1, FlashinferCascade [^12] https://github.com/flashinfer-ai/flashinfer, v0.1.4 (Ye et al., 2024a), and BatchLLM’s kernels. We compare the performance among different kernels under two scenarios: the pure decoding scenario (request count 256) and the chunked-prefill scenario where a token-batch includes both prefill chunks and decoding tokens (7 prefill requests and 256 decoding requests). Besides that, how many requests one single prefix-group have, and various length setting of both shared prefix and non-shared tokens, are non-trivial. Our experiments are conducted based on these insights.
According to the results, we can see that the Triton-based FlashAttention still has a performance gap to the CUDA-based FlashAttention, although we have already tuned the hyparameters using Triton’s autotuner. (Note we omit the performance under chunked-prefill setting because it’s even worse.) This indicates the high-level Triton compiler still have a room to achieve the performance of low-level CUDA compiler for complex computations (for example, we observe more shared memory bank conflict of the Triton FlashAttention than the CUDA version).
@@ 6.3.3. Kernel Performance Comparison
Generally, our kernel performs competitively, while some bad cases cannot be ignored. Take the scenaro where there are multiple requests with the same sharing prefix, and the common part is much longer than the distinct part, as an example. For these long-key-value cases, split-k could be a solution through which the calculation units on GPUs could be fully utilized. We’ll still focus on how to make optimization in the future.
6.3.4. Global Prefix Identification Overhead
We have measured the time of the global prefix identification described in Sec.4.2. The time range of the overhead starts before adding the token ids of each prompt to form the basic prefix tree and ends after generating the final list of prefix-sharing groups. As for the industry workload in Sec.6.2.2, the total overhead of the 8000 requests is 1.51 seconds. While the time to process the 8000 requests with the input of prefix-sharing groups format is 919.54 seconds. The global prefix identification procedure nearly has no overhead compared with the end-to-end execution time.
7. Related Work

### v3

@@ 3.4 End-to-end System Integration
Figure 6: Microbenchmark evaluation. The setting m/n (like 2000/200) indicates the length of shared prefix/non-shared context, sd means sharing degree. For subfigure (e), lengths are perturbed around base values using a Poisson distribution. The vLLM setting with ’+ p’ (’+ c’) means prefix-caching (chunked-prefill) enabled.
4 Evaluation
4.1 Setup
@@
Baseline Specification. We compare BatchLLM with vLLM and SGLang, two state-of-the-art LLM inference engines with prefix sharing, for the end-to-end comparison to demonstrate the efficacy (Sec.4.2). We integrate all our design in vLLM v0.6.4 and the baseline for comparison is also vLLM v0.6.4 (except as noted).
@@
We enable the prefix-caching and chunked-prefill in vLLM baseline for the end-to-end comparison in Sec.4.2. The prefix-caching uses the implicit LRU-based caching policy. We use 2048 as the token-batch size for chunked-prefill of vLLM baseline to maximize its throughput.
@@
Hardware and System Specification. We conduct our experiments on both NVIDIA A100 80GB GPUs (both single GPU and 2-GPUs) and AMD MI200 GPU. For multi-GPU experiments on NVIDIA A100, we utilize tensor parallelism (TP) with size of 2 (denoted as A100-TP2). The CUDA version on the NVIDIA platform is 12.1. The ROCm version on the AMD platform is 6.1. The CPU is AMD EPYC 7V12 CPU @2.45GHz. The operation system is Ubuntu 20.04.
@@
Models and Dataset. We evaluate BatchLLM and the baselines with Llama-3.1-8B AI@Meta (2024), Qwen-2.5-14B Qwen et al. (2025) and Llama-3.1-70B (4-bit GPTQ), with a set of input/output with different data distributions, and Qwen-2.5-7B for a typical industry task. As for the basic evaluation, we use three different settings of the length of common prefix and distinct prompt. The length of prefix/distinct are 2000/200, 2000/1000 and 1000/1000 respectively, while the length of generated tokens is 100. We evaluate the three settings under different share-degrees. The share-degree means the number of distinct prompts sharing the same prefix. Note that all requests are randomly shuffled. The above input/output settings are very representative as they come from the real industry workloads running on thousands of GPUs. We use a uniform length to avoid analysis inconvenience caused by different lengths. Additionally, to assess performance under more dynamic lengths, we conduct experiments on Llama-3.1-8B (A100, TP=1) where the lengths of the shared prefix, distinct context, and generated tokens are perturbed around base values using a Poisson distribution. These experiments help evaluate the robustness of BatchLLM and baselines to variations in sequence lengths. The microbenchmark evaluation is in Sec.4.2.1. As for the industry tasks, we conduct the evaluation on a typical scenario in web search engines, which is the snippet generation task that generates the snippet of a document for the user query. We conduct the breakdown experiments in Sec.4.3 to demonstrate the efficacy of each technique.
4.2 End to End Comparison
@@
Figure 7: Microbenchmark evaluation of different sharing-degrees and different shared-prefix lengths on A100. Given 6400 requests. This evaluation is based on Llama-3.1-8B-Instruct, while the length of non-shared context is 200, and the generated length is 100.
4.2.1 Basic Evaluation
[…]
@@ 4.2.1 Basic Evaluation
The impact of sharing-degree and shared prefix. We also compare the throughput of different settings after changing the sharing-degree and the length of shared prefix. Fig.7 shows that the speedup becomes more significant when more requests share the same prefix and the portion of the common prefix becomes larger, which contributes to a bigger reuse ratio, and BatchLLM can save more computation and memory for this case. With lower sharing-degree and shorter common prefix (like sd is 2 and the length of prefix is 2000), the acceleration of BatchLLM is $1.7\times$$1.5\times$, comparing with the best setting of vLLM and SGLang. When the length of shared prefix is 16000, and the share degree is 16, the best performance of vLLM and SGLang are 0.49 and 0.56, while BatchLLM achieves $10.8\times$$9.5\times$ speedup.
@@
Comparison between vLLM + request sorting and BatchLLM. We perform a preprocess for the setting of vLLM + chunked-prefill + prefix-cahcing: use python.sorted() to gather all requests with the same prefix together [^4] Here the length of shared prefix/non-shared context/generated tokens are 2000/200/100, while the sharing degree is 16. The model is Llama-3.1-8B. And the GPU is NVIDIA A100.. This can help vLLM to cache the tokens better. After this preprocess, vLLM’s throughput is boosted from 6 to 13.02, while the throughput of BatchLLM is 18. One reason is that vLLM cannot simplify the attention calculation when there are more than 1 common prefix in the batch every step of its inference, while BatchLLM could handle it. Besides, BatchLLM has a better token-batching strategy than vLLM.
Comparison with the latest prefix-caching advances by early 2026. BatchLLM is implemented based on an old vLLM version, v0.6.4. The latest vLLM version has integrated many of the most recent prefix-sharing technologies from the community, and we make a comparison against it. We evaluate the industry workload in this paper with the latest vLLM (the vLLM’s github commit by 1/16/2026). Its throughput on A100 is 6.57 for this workload with its latest prefix-caching optimization, while BatchLLM achieves 8.67 even with a very old vLLM as the implementation base.
@@
Table 1: Industry workload evaluation.
@@
BatchLLM 8.67($1.30\times$) 4.24($1.26\times$)
@@
Figure 8: The length distribution of shared prefix/non-shared context in the dataset of the industry workload.
4.2.2 Industry Workload Evaluation
@@
This industry workload is similar to the snippet generation task described in the recent work Wang et al. (2023), which is also described in Sec.2.1. The original input has two levels of prefixes: the global shared instruction to extract the information from the document-query pair, and the shared document between different groups of queries. The first-level prefix is very short. With the optimization described in Sec.3.2, the first-level prefixes are enlarged (i.e., some of the second-level prefix of the document is merged into the first-level prefix). We have analyzed the token saving ratio according to Eq.1. It shows that the saving ratio are nearly the same between the multi-level and the enlarged single level prefixes, 56% for the former and 55% for the latter. Besides, the DP algorithm only introduces less than 0.01% overhead.
@@
The data distribution of this workload is shown in Fig.8. The average share-degree (i.e., the number of the distinct prompt that share the same prefix) is about 3, the average common prefix length is about 1570 tokens, and the average distinct prompt length is about 30 tokens. We sample 8000 queries for this case study.
[…]
4.2.3 Heavy-tailed decoding with varied output lengths.
We evaluate heavy-tailed decoding to demonstrate BatchLLM’s efficacy for varied output lengths, using a Pareto-distributed output length configuration.
[…]
Table 2: End-to-end throughput under the heavy-tail workload.
Configuration Throughput (req/s) Speedup
[…]
4.3 Breakdown Analysis
To better isolate the performance improvements of BatchLLM, we provide the breakdown analysis to evaluate the individual effects of our core techniques: prefix grouping, request reordering (token-batching), and kernel fusion.
4.3.1 Effect of Prefix Grouping
Algorithm Effectiveness on Multi-level Prefix Inputs. To demonstrate the effectiveness of our proposed prefix identification algorithm, we conduct an ablation study comparing it against a baseline sorting-based heuristic approach for multi-level prefix inputs. The heuristic method relies on lexicographical sorting and greedy scanning, grouping requests only if their longest common prefix exceeds a rigid 50% length threshold relative to the shortest member. In contrast, our approach automatically discovers shared prefixes of arbitrary lengths without rigid ratio constraints, merging paths based on a dynamic gain formula.
We construct a synthetic dataset to evaluate edge cases where prompts share significant context but fall short of heuristic thresholds. Each prompt consists of $1000$ tokens formulated as: group_prefixgroup\_prefix + sub_prefixsub\_prefix + random_suffixrandom\_suffix. We generate $6400$ prompts divided into $50$ global groups. Each group contains $64$ subcategories, with $2$ prompts per subcategory. We test two configurations: Setting A uses a $490$-token group_prefixgroup\_prefix and an $11$-token sub_prefixsub\_prefix, while Setting B uses a $400$-token group_prefixgroup\_prefix and a $101$-token sub_prefixsub\_prefix.
In both settings, the prefix ratio within the same subcategory is $50\%$ (e.g., $(490+11)/1000=50.1\%$), but the prefix ratio across different subcategories within the same global group falls strictly below $50\%$ (e.g., $490/1000=49\%$).
Table 3: Performance comparison of prefix grouping algorithms across $6400$ requests. The baseline executes without context sharing (CS) or BatchLLM optimizations.
Configuration Saving Ratio Throughput (req/s)
Setting A (group_prefix=490group\_prefix=490)
Baseline (No CS) 0.0% 10.7
Heuristic CSGroup $\sim$25.0% 12.4
BatchLLM 48.6% 15.9
Setting B (group_prefix=400group\_prefix=400)
Baseline (No CS) 0.0% 10.7
Heuristic CSGroup $\sim$25.0% 12.3
BatchLLM 39.7% 14.5
As shown in Table 3, the heuristic approach is bottlenecked by its 50% threshold. It successfully groups prompts within the same subcategory (forming $3200$ groups of $2$ members), yielding a limited token saving ratio of $\sim$25% and throughput of $\sim$12.4 prompts/s.
[…]
@@ 4.3.1 Effect of Prefix Grouping
KV Cache Reusing Effect. To show that this grouping optimizes the KV reuse, we evaluate the throughput and the token saving ratio in both microbenchmarks and the industry task. Note that all samples in the the industry task are not sorted too.
@@
Figure 9: The token saving (or reusing) ratio of different sharing-degrees and different shared-prefix lengths on A100, based on the microbenchmark. This evaluation is based on Llama-3.1-8B-Instruct, while the length of non-shared context is 200, and the generated length is 100.
@@
As shown in Fig.9, we can see the token saving ratio is improved in BatchLLM across different settings. And it becomes higher when the sharing-degree is larger. Specifically, in the microbenchmark setting where length of shared prefix is 2000 and the sharing-degree is 16, BatchLLM achieves 85.2% token saving ratio, the upper bound in this scenario, comparing to the 40.1% token saving ratio in vLLM + chunked-prefill + prefix-caching. With the increase of the requests number and the length of prefix, it can be expected that the implicit LRU cache policy of vLLM’s prefix-caching can suffer more from the eviction of the reusable KV context, while the inference could benefit more from BatchLLM. For example, when the length of shared prefix is 16000, and the share degree is 16, the token reusing ratio of vLLM and SGLang are 6.3% and 5.2%, while the reusing ratio of BatchLLM is 92.6 %, much better than the other two baselines.
Global Prefix Identification Overhead. We have measured the time of the global prefix identification described in Sec.3.2. The time range of the overhead starts before adding the token ids of each prompt to form the basic prefix tree and ends after generating the final list of prefix-sharing groups. As for the industry workload in Sec.4.2.2, the total overhead of the 8000 requests is 1.51 seconds. While the time to process the 8000 requests with the input of prefix-sharing groups format is 919.54 seconds. The global prefix identification procedure nearly has no overhead compared with the end-to-end execution time.
4.3.2 Effect of Request Reordering and Token-Batching
@@
Fig.10 shows the number of per-iteration tokens, using the same workload with Fig.2. It clearly shows that the request reordering and token-batching optimization in BatchLLM successfully reduces the "valleys" of the iterations, thus increasing the overall GPU utilization.
@@
Figure 10: Token number per iteration with token-batching optimization (Sec.3.3), showing significant valley reduction compared to Fig.2’s baseline. Prefix-sharing disabled to demonstrate general batched/offline inference efficacy.
Table 4: Token-Batching Breakdown.
@@
4.3.3 Effect of Horizontal Kernel Fusion
The existing works use separate kernels to compute the partial Attention on the prefix and the distinct KV context. This will lead to the tail effect of each kernel and the launch overhead of these kernels. BatchLLM horizontally fuses the two partial Attention calculation into one kernel. Different parts of the computation are performed in different thread blocks. Note that the problem shape of the two parts are different. BatchLLM applies different tiling configuration for the two parts to achieve the best performance. It uses the common auto-tuning method to find the best configuration.
We implement the Attention in BatchLLM through OpenAI Triton language Tillet et al. (2019). We use the builtin autotuner of Triton to tune the tiling size of the Attention implementation in BatchLLM. Besides that, some ahead-of-time compilation methods are applied to reduce the launching overhead, including warm-up process on NVIDIA GPUs, and AOTriton [^5] https://github.com/ROCm/aotriton on AMD GPUs.
@@ 4.3.3 Effect of Horizontal Kernel Fusion
Figure 11: The performance comparison between the baseline, CascadeInfer and BatchLLM. The setting i/j/k (like 256/2048/all) means the length of global shared prefix, the length of distinct parts for each request, and how many requests in one prefix-sharing group.
We compare BatchLLM’s prefix-shared Attention implementation with several baselines on both NVIDIA A100 GPU and AMD MI200 GPU. Fig.11 shows the performance comparison of different implementations, including Triton official FlashAttention implementation, official FlashAttention v2.6.1, FlashinferCascade v0.1.4 Ye et al. (2024b), and BatchLLM’s kernels. We compare the performance among different kernels under two scenarios: the pure decoding scenario (request count 256) and the chunked-prefill scenario where a token-batch includes both prefill chunks and decoding tokens (7 prefill requests and 256 decoding requests).
Results show that the Triton-based FlashAttention still has a performance gap to the CUDA-based FlashAttention, although we have already tuned the hyparameters using Triton’s autotuner. This indicates the high-level Triton compiler still have a room to achieve the performance of low-level CUDA compiler for complex computations.
@@
4.4 Limitations and Deployment Considerations
While BatchLLM demonstrates superior performance in the industrial scenarios detailed in Sec.2.1, it is subject to specific limitations. Primarily, the system’s design may adversely affect individual request latency, rendering it less suitable for latency-critical workloads. A potential mitigation is the development of a hybrid architecture that dynamically toggles between streaming and batching modes. Specifically, when the system encounters a backlog of tasks—transitioning into a batching-intensive regime—and identifies a high global prefix sharing ratio, it could employ BatchLLM techniques to maximize throughput and reduce queue depth. Otherwise, the system should utilize standard execution paths to prioritize low latency. We leave the implementation of such a hybrid system for future work. Secondly, the advantages of BatchLLM are diminished in environments lacking common prefixes. In the absence of prefix caching, performance gains are restricted solely to those provided by token batching. Therefore, we recommend deploying BatchLLM specifically for prefix-heavy, throughput-oriented workloads rather than general-purpose tasks.
5 Related Work